# SonicSim: A customizable simulation platform for speech processing in moving sound source scenarios

### v1

@@ 4.2 Benchmark models
We selected several popular models that have demonstrated excellent performance in speech separation as benchmarks, including Conv-TasNet (Luo & Mesgarani, 2019), DPRNN (Luo et al., 2020), DPTNet (Chen et al., 2020a), SuDORM-RF (Tzinis et al., 2020), A-FRCNN (Hu et al., 2021), SKIM (Li et al., 2022d), TDANet (Li et al., 2023), BSRNN (Luo & Yu, 2023b), TF-GridNet (Wang et al., 2023), Mossformer (Zhao & Ma, 2023), and Mossformer2 (Zhao et al., 2024). For detailed information about the benchmark models, see Appendix C.2. The training object of the models is available in Appendix C.3. The hyperparameter configuration and all training settings are available in Appendix C.4. The PyTorch implementations and pre-trained weights for all benchmark models are publicly available444https://github.com/JusperLee/SonicSim/tree/main/separation.
@@ 4.3 Dataset details
We report the performance of trained baseline models using the LRS2-2Mix (Li et al., 2023), Libri2Mix (Cosentino et al., 2020), and SonicSet datasets, which we tested on a real-world moving speech separation dataset collected by us. We used the test sets from LRS2-2Mix and Libri2Mix to test the models’ generalization and transfer performance across different scenarios. Please refer to Appendix C.5 for detailed descriptions of these datasets.
@@ 4.4 Evaluation metrics
We employed a series of evaluation metrics to assess the performance of the speech separation benchmark models comprehensively. These metrics cover multiple dimensions of audio quality, including signal quality (SI-SNR (Le Roux et al., 2019) and SDR (Vincent et al., 2006)), speech intelligibility (STOI (Taal et al., 2011) and WER), and subjective quality perception (PESQ (Rix et al., 2001)). Details are given in Appendix C.6. For these metrics, higher values are preferable for all indicators except for WER, where a lower value is better. From the perspective of practical applications, WER, a speech recognition metric, is the most important metric because speech separation is usually a preprocessing step for speech recognition.
@@ 4.5 Results and analysis
Comparison on real-recorded datasets. To validate the acoustic simulation gap between synthetic datasets and real data, we constructed a real-world moving sound source dataset consisting of 5 hours of recorded audio (details in Appendix C.5), encompassing various complex acoustic environments and different types of background noise (environmental and musical noise). In Tables 2 and 3, we trained models using the LRS2-2Mix and Libri2Mix datasets to compare their performance with models trained on the SonicSet dataset. The LRS2-2Mix dataset (Li et al., 2023) contains real-world noise and reverberation. The Libri2Mix dataset (Cosentino et al., 2020) is currently the largest synthetic speech separation dataset, and its data scale is closest to SonicSet among public speech separation datasets. We trained different models on these datasets and tested them on the real-world recorded datasets. The results demonstrated that models trained on the SonicSet dataset achieved overall the best separation performance on the real-world recorded datasets. This finding indicates that the SonicSet dataset excels in simulating more realistic acoustic scenes, making it effective for model training and evaluation.
Comparison on the SonicSet dataset. The results of the speech separation are shown in Table 4. Different models exhibit varying performance under environmental and musical noise conditions. Among the RNN-based models, DPRNN and BSRNN performed relatively poorly when handling musical noise. In contrast, CNN-based models (SuDoRM-RF and A-FRCNN) demonstrated more balanced performance across both environmental and musical noise scenarios. Among Transformer-based models, Mossformer2 showed the SOTA performance, especially in handling complex noise. This superiority can be attributed to incorporating the multi-head attention mechanism, which effectively captures long-term dependencies and complex frequency variations (Vaswani et al., 2017).
@@ 5.1 Problem definition
The speech enhancement task aims to extract high-quality target speech from a noisy signal, reducing or eliminating background noise, as shown in Figure 4(b). This task is crucial in applications such as speech recognition, speech communication, and hearing aids. A detailed explanation of this task pipeline is available in Appendix D.1.
@@ 5.2 Benchmark models
In the speech enhancement experiments, we selected several popular models: DCCRN (Hu et al., 2020), Fullband (Hao et al., 2021), FullSubNet (Hao et al., 2021), Fast-FullSubNet (Hao & Li, 2022), FullSubNet+ (Chen et al., 2022b), TaylorSENet (Li et al., 2022a), GaGNet (Li et al., 2022c), G2Net (Li et al., 2022b) and Inter-SubNet (Chen et al., 2023). The details of the models are available in Appendix D.2. The training object of the model is available in Appendix D.3. The hyperparameter configuration and all training settings are available in Appendix D.4.
@@ 5.3 Evaluation metrics
In the speech enhancement task, we use a series of evaluation metrics to comprehensively evaluate the performance of the speech separation baseline model, including signal quality (SI-SNR (Le Roux et al., 2019) and SDR (Vincent et al., 2006)), speech intelligibility (STOI (Taal et al., 2011) and CER), and subjective quality perception (PESQ (Rix et al., 2001), DNSMOS (Reddy et al., 2021) and SigMOS (Ristea et al., 2024)). Since the real-world dataset, RealMAN is a Chinese dataset, we use CER as the evaluation metric for speech recognition. For these metrics, except for CER, the higher the value, the better, while for CER, the lower the value, the better. From an application perspective, CER is the most important metric, as speech enhancement is usually a pre-processing step for speech recognition. The PyTorch implementations and pre-trained weights for all baseline models have been made publicly available555https://github.com/JusperLee/SonicSim/tree/main/enhancement.
@@ 5.5 Results and analysis
Comparison on real-recorded dataset. We utilized the real-world recorded moving source dataset, RealMAN (test set), to evaluate trained speech enhancement models on various speech enhancement datasets, including VoiceBank-DEMAND, DNS Challenge, and SonicSet. VoiceBank-DEMAND is a smaller dataset, while the DNS Challenge dataset comprises approximately 700 hours of data, with a data scale comparable to that of SonicSet. We employed the DNSMOS (Reddy et al., 2021) tool to calculate subjective metrics, while Character Error Rate (CER) results were evaluated using the same speech recognition model666https://huggingface.co/espnet/pengcheng_guo_wenetspeech_asr_train_asr_raw_zh_char as used with RealMAN. The results presented in Table 5 indicated that models trained on the SonicSet dataset achieved overall the best results across multiple metrics. Please note that for most practical applications, CER is the most important metric. Based on these findings, we infer that the synthetic dataset SonicSet effectively simulates the acoustic environments of real moving source scenarios. This provides significant flexibility and cost-effectiveness for constructing large-scale datasets, positioning synthetic datasets as an effective alternative for addressing complex environments.
Comparison on the SonicSet dataset. By comparing the performance of different speech enhancement models on the SonicSet dataset, we analyzed their enhancement effects under various types of noise. Table 6 presents the performance of the benchmark models. Sub-band splitting methods (such as FullSubNet, Fast-FullSubNet, FullSubNet+, and Inter-SubNet) performed well across multiple metrics. Their interaction processing mechanism between sub-bands significantly improved the models’ ability to handle frequency band dependencies, resulting in better speech clarity and noise suppression. In contrast, although TaylorSENet and GaGNet showed relatively similar performance on SI-SNR and SDR, they fell slightly behind in the MOS series scores.
@@ 6 Conclusions
Appendix A Dataset analysis
@@
Appendix B SonicSet demo
@@
Appendix C Benchmark I: speech separation
C.1 Problem definition
@@ C.2 Benchmark models
TDANet (Li et al., 2023): Simplifies redundant connections in A-FRCNN and introduces a top-down attention mechanism to selectively extract multi-level acoustic features, further improving processing efficiency. The introduction of attention module further enhances the model’s ability to extract complex correlations of temporal information at different scales.
@@
BSRNN (Luo & Yu, 2023b): Proposes a frequency-band splitting method, modeling both within frequency bands and across the time dimension using BLSTM, marking the first application of band-splitting methods in speech separation. This unique approach allows for more granular control over the frequency components, with better performance in tasks that rely more on fine frequency features such as music separation tasks.
@@
TF-GridNet (Wang et al., 2023): Extends DPRNN by modeling in both the time and frequency dimensions and integrates information from both dimensions using a multi-head attention module, leading to more refined and comprehensive speech separation. Compared to previous attention based speech separation works, this work models in both time and frequency domains, and reduces the length load of the attention module, making it more conducive to capturing contextual relationships.
@@
Mossformer (Zhao & Ma, 2023): Combines convolution and self-attention in a gated, single-head Transformer architecture, utilizing full self-attention within local blocks and a linearized, low-cost self-attention mechanism for global modeling, significantly improving separation performance. The model also utilizes an attention gathering mechanism with simplified single head self attention. The model achieved good performance by integrating multiple modules.
@@
Mossformer2 (Zhao et al., 2024): Enhances Mossformer by introducing a feed-forward sequential memory network (FSMN) recurrent module, improving long-term dependency modeling and further boosting speech separation effectiveness. This FSMN blocks are mainly enhanced by using gated convolution units (GCUs) and dense connections. In addition, bottleneck layers and output layers have been added to control the flow of information. Based on the above content, the model integrates the loop module into the MossFormer framework, providing the ability to model remote, coarse scale dependencies, and fine scale loop patterns.
C.3 Training object
@@
C.4 Implementation details
During training, since all model hyperparameters were originally configured for the WSJ0-2Mix dataset (Hershey et al., 2016) and suited for audio with an 8 kHz sampling rate, we doubled the window length and window shift of the encoder and decoder to adapt to the 16 kHz sampling rate datasets. Aside from this adjustment, all other hyperparameters remained unchanged. For training, we randomly sampled 3-second audio clips from the training set. The batch size was set to 1, and the Adam optimizer (Kingma, 2014) was used with an initial learning rate of . The learning rate was halved whenever the validation loss did not decrease for five consecutive epochs. We applied gradient clipping to limit the maximum L2 norm of the gradients to 5. The training ran for a maximum of 500 epochs, with early stopping applied if the validation loss did not improve for 10 consecutive epochs.
For each benchmark model, we selected the best model based on its performance on the validation set for testing. In the SonicSet evaluation, we used the full test set and employed a 6-second inference window with a 3-second sliding window to process long audio sequences. For the LRS2-2Mix dataset, we used 2-second audio segments for both training and testing. For the Libri2Mix dataset, we trained and tested the models using 3-second audio segments. All experiments were conducted on a server equipped with 8 NVIDIA 4090 GPUs.
C.5 Real-world dataset details
In the speech separation experiments, we used SonicSet, public datasets, and real-world datasets. Each mixed audio contained two different speakers, with a sampling rate of 16 kHz. The baseline models pre-trained on the SonicSet training set were trained using a dynamic augmentation strategy. Specifically, during training, we randomly selected a set of data, and two speakers were randomly chosen from three available audio tracks for mixing. Depending on the task requirements, either environmental noise or musical noise was then mixed accordingly. Finally, a 3-second segment was randomly extracted from the 60-second audio for model training.
To evaluate the gap between the SonicSet dataset and real-world conditions, as well as the model’s transferability to real-world scenarios, we collected a small-scale real-world moving sound source dataset and repeated the relevant experiments. First, we randomly selected 30 audio samples from 10 scenes in the SonicSet validation set, totaling 5 hours of audio. All data was clean, without reverberation or noise. During recording, one participant used the 2023 MacBook Pro’s speakers to play audio while moving around randomly within the same scene to obtain the ground-truth audio. In addition, we used the same steps to play environmental and music noise from fixed positions to obtain noise data. The audio and noise were recorded using a Logitech Blue Yeti Nano omnidirectional microphone fixed in position, with a recording sample rate of 16 kHz and 32-bit depth. After recording, we clipped the audio and noise based on the recorded start and stop positions to ensure alignment with the original files. We then recorded another real-world dataset in 10 similar real-world scenes using the corresponding original audio. Finally, we mixed the data using the same method as SonicSet to construct a dataset for testing model performance. The layout of the real-world scenes is illustrated in Figure 6.
C.6 Evaluation metrics’ details

### v2

@@ 4.2 Benchmark models
We selected several popular models that have demonstrated excellent performance in speech separation as benchmarks, including Conv-TasNet (2019), DPRNN (2020), DPTNet (2020a), SuDORM-RF (2020), A-FRCNN (2021), SKIM (2022d), TDANet (2023), BSRNN (2023b), TF-GridNet (2023), Mossformer (2023), and Mossformer2 (2024). For detailed information about the benchmark models, see Appendix D.2. The PyTorch implementations and pre-trained weights for all benchmark models are publicly available555https://github.com/JusperLee/SonicSim/tree/main/separation.
@@ 4.3 Dataset details
We report the performance of trained baseline models using the LRS2-2Mix (Li et al., 2023), Libri2Mix (Cosentino et al., 2020), and SonicSet datasets, which we tested on a real-world moving speech separation dataset collected by us. We used the test sets from LRS2-2Mix and Libri2Mix to test the models’ generalization and transfer performance across different scenarios. Please refer to Appendix D.3 for detailed descriptions of these datasets.
4.4 Implementation details
To adapt the WSJ0-2Mix model (originally for 8 kHz) to our 16 kHz dataset, we doubled the window length and step size of the encoder and decoder, keeping other hyperparameters unchanged. We trained the model using the Adam optimizer (Kingma, 2014) with a batch size of 1 on randomly sampled 3s audio clips, employing negative SNR as the loss function. The initial learning rate was set to , with gradient clipping and early stopping strategies applied. Detailed hyperparameters and training settings are provided in Appendix D.4 and D.5. For experiments, we used SonicSet, public datasets, and real-world datasets, with each mixed audio containing two different speakers at 16 kHz. The baseline model was pre-trained on SonicSet using a dynamic augmentation strategy (Luo & Yu, 2023b): two speakers are randomly selected from three audio tracks for mixing, adding environmental or music noise as required, and a 3s segment is randomly extracted from the 60s audio for training. In SonicSet, we used the full test set and processed with a 6s inference window and a 3s sliding window. For the LRS2-2Mix (Li et al., 2023) and Libri2Mix (Cosentino et al., 2020) datasets, 2s and 3s audio clips were used for training and testing, respectively.
@@ 4.5 Evaluation metrics
We employed a series of metrics to assess the performance of the speech separation benchmark models comprehensively. These metrics cover multiple dimensions of audio quality, including signal quality (SI-SNR (Le Roux et al., 2019) and SDR (Vincent et al., 2006)), speech intelligibility (STOI (Taal et al., 2011) and WER), and subjective quality perception (PESQ (Rix et al., 2001), NISQA (Mittag et al., 2021) and SigMOS (Ristea et al., 2024) ). Details are given in Appendix D.6.
@@ 4.6 Results and analysis
Comparison on real-world datasets. To validate the acoustic simulation gap between synthetic datasets and real data, we constructed a real-world moving sound source dataset consisting of 5 hours of recorded audio (details in Appendix D.3), namely RealSEP, encompassing various complex acoustic environments and different types of background noise (environmental and musical noise). In Tables 2 and 3, we trained models using the LRS2-2Mix and Libri2Mix datasets to compare their performance with models trained on the SonicSet dataset. The LRS2-2Mix dataset (Li et al., 2023) contains real-world noise and reverberation. The Libri2Mix dataset (Cosentino et al., 2020) is currently the largest synthetic speech separation dataset, and its data scale is closest to SonicSet among public speech separation datasets. We trained different models on these datasets and tested them on the RealSEP datasets. In addition, we supplemented the experimental results by recording audio data of direct conversations between individuals (HumanSEP, details in Appendix D.3). We then evaluated the NISQA and WER results of the speech separation model trained on different data sets, as shown in Table 10 in the Appendix. The results demonstrated that models trained on the SonicSet dataset achieved overall the best separation performance on the RealSEP and HumanSEP datasets. This finding indicates that the SonicSet dataset excels in simulating more realistic acoustic scenes, making it effective for model training and evaluation.
Comparison on the SonicSet dataset. The results of the speech separation are shown in Table 4. Different models exhibit varying performance under environmental and musical noise conditions. Among the RNN-based models, DPRNN and BSRNN performed relatively poorly when handling musical noise. In contrast, CNN-based models (SuDoRM-RF and A-FRCNN) demonstrated more balanced performance across both environmental and musical noise scenarios. Among Transformer-based models, Mossformer2 showed the SOTA performance, especially in handling complex noise. This superiority can be attributed to incorporating the multi-head attention mechanism, which effectively captures long-term dependencies and complex frequency variations (Vaswani et al., 2017). The detailed discussion is in Appendix D.7.
@@ 5.1 Problem definition
The speech enhancement task aims to extract high-quality target speech from a noisy signal, reducing or eliminating background noise, as shown in Figure 4(b). This task is crucial in applications such as speech recognition, speech communication, and hearing aids. A detailed explanation of this task pipeline is available in Appendix E.1.
@@ 5.2 Benchmark models
5.3 Implementation details
The hyperparameter settings for all baseline models remain consistent with the original papers. We trained the models using the Adam optimizer (Kingma, 2014) with an initial learning rate of 0.001. If the validation loss does not decrease for five consecutive epochs, we reduce the learning rate by a factor of 0.5. All audio samples are processed at a 16 kHz sampling rate. Training employs early stopping and halting if the validation loss doesn’t improve for 10 epochs. During testing, we use the same inference strategy as in the speech separation task to handle long audio samples. For the speech enhancement experiments, we trained models on the VoiceBank-DEMAND (Valentini-Botinhao et al., 2016), DNS Challenge (Reddy et al., 2020), and SonicSet datasets using single-channel speech signals sampled at 16 kHz. Baseline models were pre-trained on SonicSet with the same dynamic enhancement strategy as in separation model training, extracting 3s segments from 60s audio for training. The training objective is detailed in Appendix E.3. For SonicSet, we used the full test set with a 6s inference window and a 3s sliding window to process long audio sequences. On the VoiceBank-DEMAND and DNS Challenge datasets, 2s and 3s audio clips were used for training and testing, respectively.
@@ 5.4 Evaluation metrics
In the speech enhancement task, we use a series of evaluation metrics to comprehensively evaluate the performance of the speech separation baseline model, including signal quality (SI-SNR (Le Roux et al., 2019) and SDR (Vincent et al., 2006)), speech intelligibility (STOI (Taal et al., 2011) and CER), and subjective quality perception (PESQ (Rix et al., 2001), NISQA (Mittag et al., 2021), DNSMOS (Reddy et al., 2021) and SigMOS (Ristea et al., 2024)). Since the real-world dataset, RealMAN is a Chinese dataset, we use CER as the evaluation metric for speech recognition. For these metrics, except for CER, the higher the value, the better, while for CER, the lower the value, the better. From an application perspective, CER is the most important metric, as speech enhancement is usually a pre-processing step for speech recognition. The PyTorch implementations and pre-trained weights for all baseline models have been made publicly available666https://anonymous.4open.science/r/SonicSim-ICLR2025/enhancement.
@@ 5.6 Results and analysis
Comparison on real-recorded dataset. We utilized the real-world recorded moving source dataset, RealMAN (test set), to evaluate trained speech enhancement models on various speech enhancement datasets, including VoiceBank-DEMAND, DNS Challenge, and SonicSet. VoiceBank-DEMAND is a smaller dataset, while the DNS Challenge dataset comprises approximately 700 hours of data, with a data scale comparable to that of SonicSet. We employed the DNSMOS (Reddy et al., 2021) tool to calculate subjective metrics, while Character Error Rate (CER) results were evaluated using the same speech recognition model777https://huggingface.co/espnet/pengcheng_guo_wenetspeech_asr_train_asr_raw_zh_char as used with RealMAN. The results presented in Table 5 indicated that models trained on the SonicSet dataset achieved overall the best results across multiple metrics. Please note that for most practical applications, CER is the most important metric. Moreover, we supplemented the experimental results by constructing a dataset called HumanENH (details in Appendix E.4) by recording human speech audio and noise data. We then evaluated the NISQA and WER results of the speech enhancement models trained on different datasets, as shown in Table 11 in the appendix.
Based on these findings, we infer that the synthetic dataset SonicSet effectively simulates the acoustic environments of real moving source scenarios. This provides significant flexibility and cost-effectiveness for constructing large-scale datasets, positioning synthetic datasets as an effective alternative for addressing complex environments.
Comparison on the SonicSet dataset. By comparing the performance of different speech enhancement models on the SonicSet dataset, we analyzed their enhancement effects under various types of noise. Table 6 presents the performance of the benchmark models. Sub-band splitting methods (such as FullSubNet, Fast-FullSubNet, FullSubNet+, and Inter-SubNet) performed well across multiple metrics. Their interaction processing mechanism between sub-bands significantly improved the models’ ability to handle frequency band dependencies, resulting in better speech clarity and noise suppression. In contrast, although TaylorSENet and GaGNet showed relatively similar performance on SI-SNR and SDR, they fell slightly behind in the MOS series scores. The detailed discussion is in Appendix E.5.
@@ 6 Conclusions
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China (No. 2021ZD0200301) and the National Natural Science Foundation of China (No. U2341228).
Appendix A Environmental factor
We randomly selected 300 audio files from the Librispeech (Panayotov et al., 2015) test-clean set. 200 of these were used to generate noisy mixtures containing two speakers, and the other 100 were used to generate noisy audio from a single speaker. The same audio files were used in both cases. To fix the scene, we chose a cuboid room with no obstructions but included room materials and fixed the positions of the microphone, sound source, and noise. In the experiment on occlusion, we used Habitat-sim’s API “add_object” to place objects between the microphone and the sound source to simulate the occlusion in sound wave propagation. When studying the effect of room shape, we chose a cylindrical room from the Matterport3D dataset (Chang et al., 2017) and fixed the relative positions of the microphone, sound source, and noise. In addition, we completely removed the root material information from the experiment to explore the effect of the material.
We selected the two methods with the best speech separation and enhancement performance and verified the impact of different environmental factors on the model performance. The experimental results are shown in Tables 7 and 8. The results show that material has the most significant effect on model performance, while the impact of room shape is relatively small. However, all these factors have a certain degree of interference in model performance. Therefore, a simulated dataset using these environmental factors is generated to train the model and improve its generalization ability.
| Methods | with/without obstacles | cylinder/cuboid | with/without materials | |||
| SDR (↑) | WER (%) (↓) | SDR (↑) | WER (%) (↓) | SDR (↑) | WER (%) (↓) | |
| Mossformer | 10.70/11.84 | 29.90/24.31 | 11.18/11.84 | 27.17/24.31 | 11.84/10.12 | 24.31/31.10 |
| Mossformer2 | 10.41/11.60 | 31.89/25.86 | 10.78/11.60 | 30.54/25.86 | 11.60/9.87 | 25.86/33.58 |
| with/without obstacles | cylinder/cuboid | with/without materials | ||||
| Methods | SDR (↑) | WER (%) (↓) | SDR (↑) | WER (%) (↓) | SDR (↑) | WER (%) (↓) |
| FullSubNet | 9.67/11.05 | 25.76/19.80 | 10.56/11.05 | 21.94/19.80 | 11.05/10.23 | 19.80/21.17 |
| Inter-SubNet | 10.78/12.69 | 22.96/17.68 | 11.50/12.69 | 19.30/17.68 | 12.69/11.26 | 17.68/19.88 |
Appendix B Dataset analysis
@@
Appendix C SonicSet demo
@@
Appendix D Benchmark I: speech separation
D.1 Problem definition
@@ D.2 Benchmark models
TDANet (Li et al., 2023): Simplifies redundant connections in A-FRCNN and introduces a top-down attention mechanism to selectively extract multi-level acoustic features, further improving processing efficiency. The introduction of the attention module further enhances the model’s ability to extract complex correlations of temporal information at different scales.
@@
BSRNN (Luo & Yu, 2023b): Proposes a frequency-band splitting method, modeling both within frequency bands and across the time dimension using BLSTM, marking the first application of band-splitting methods in speech separation. This unique approach allows for more granular control over the frequency components, with better performance in tasks that rely more on fine frequency features, such as music separation tasks.
@@
TF-GridNet (Wang et al., 2023): Extends DPRNN by modeling in both the time and frequency dimensions and integrates information from both dimensions using a multi-head attention module, leading to more refined and comprehensive speech separation. Compared to previous attention-based speech separation works, this work models in both time and frequency domains and reduces the length load of the attention module, making it more conducive to capturing contextual relationships.
@@
Mossformer (Zhao & Ma, 2023): Combines convolution and self-attention in a gated, single-head Transformer architecture, utilizing full self-attention within local blocks and a linearized, low-cost self-attention mechanism for global modeling, significantly improving separation performance. The model also utilizes an attention-gathering mechanism with simplified single head self-attention. The model achieved good performance by integrating multiple modules.
@@
Mossformer2 (Zhao et al., 2024): Enhances Mossformer by introducing a feed-forward sequential memory network (FSMN) recurrent module, improving long-term dependency modeling and further boosting speech separation effectiveness. These FSMN blocks are mainly enhanced by using gated convolution units (GCUs) and dense connections. In addition, bottleneck layers and output layers have been added to control the flow of information. Based on the above content, the model integrates the loop module into the MossFormer framework, providing the ability to model remote, coarse-scale dependencies and fine-scale loop patterns.
D.3 Real-world dataset details
To evaluate the gap between the SonicSet dataset and real-world conditions, as well as the model’s transferability to real-world scenarios, we collected a small-scale real-world moving sound source dataset, namely RealSEP and repeated the relevant experiments. First, we randomly selected 30 audio samples from 10 scenes in the SonicSet validation set, totaling 5 hours of audio. All data was clean, without reverberation or noise. During recording, one participant used the 2023 MacBook Pro’s speakers to play audio while moving around randomly within the same scene to obtain the ground-truth audio. In addition, we used the same steps to play environmental and music noise from fixed positions to obtain noise data. The audio and noise were recorded using a Logitech Blue Yeti Nano omnidirectional microphone fixed in position, with a recording sample rate of 16 kHz and 32-bit depth. After recording, we clipped the audio and noise based on the recorded start and stop positions to ensure alignment with the original files. We then recorded another real-world dataset in 10 similar real-world scenes using the corresponding original audio. Finally, we mixed the data using the same method as SonicSet to construct a dataset for testing model performance. The layout of the real-world scenes is illustrated in Figure 6.
In addition, we have constructed a dialogue speech dataset called HumanSEP, which contains recordings of 16 speakers in 6 different scenarios. Specifically, in each scenario, two speakers each randomly read aloud 15 sentences selected from the LibriSpeech dataset (Panayotov et al., 2015), with an average length of 4–6 seconds. To simulate real-world usage scenarios, the speech data was randomly interspersed with environmental noise. All audio was recorded using a fixed-position Logitech Blue Yeti Nano omnidirectional microphone with recording parameters set to a 16 kHz sampling rate and 32-bit depth. The dataset was also designed to include varying degrees of overlap between speakers to increase its diversity and challenge.
D.4 Training object
@@
D.5 Implementation details
During training, since all model hyperparameters were originally configured for the WSJ0-2Mix dataset (Hershey et al., 2016) and suited for audio with an 8 kHz sampling rate, we doubled the window length and window shift of the encoder and decoder to adapt to the 16 kHz sampling rate datasets. Aside from this adjustment, all other hyperparameters remained unchanged. For training, we randomly sampled 3s audio clips from the training set. The batch size was set to 1, and the Adam optimizer (Kingma, 2014) was used with an initial learning rate of . The learning rate was halved whenever the validation loss did not decrease for five consecutive epochs. We applied gradient clipping to limit the maximum L2 norm of the gradients to 5. The training ran for a maximum of 500 epochs, with early stopping applied if the validation loss did not improve for 10 consecutive epochs.
In the speech separation experiments, we used SonicSet, public datasets, and real-world datasets. Each mixed audio contained two different speakers, with a sampling rate of 16 kHz. The baseline models pre-trained on the SonicSet training set were trained using a dynamic augmentation strategy. Specifically, during training, we randomly selected a set of data, and two speakers were randomly chosen from three available audio tracks for mixing. Depending on the task requirements, either environmental noise or musical noise was then mixed accordingly. Finally, a 3s segment was randomly extracted from the 60s audio for model training.
For each benchmark model, we selected the best model based on its performance on the validation set for testing. In the SonicSet evaluation, we used the full test set and employed a 6s inference window with a 3s sliding window to process long audio sequences. For the LRS2-2Mix dataset, we used 2s audio segments for both training and testing. For the Libri2Mix dataset, we trained and tested the models using 3s audio segments. All experiments were conducted on a server equipped with 8 NVIDIA 4090 GPUs.
D.6 Evaluation metrics’ details